# Agentic Robot: A Brain-Inspired Framework for Vision-Language-Action Models in Embodied Agents

### v1

@@ Abstract
Long-horizon robotic manipulation poses significant challenges for autonomous systems, requiring extended reasoning, precise execution, and robust error recovery across complex sequential tasks. Current approaches, whether based on static planning or end-to-end visuomotor policies, suffer from error accumulation and lack effective verification mechanisms during execution, limiting their reliability in real-world scenarios. We present Agentic Robot, a brain-inspired framework that addresses these limitations through Standardized Action Procedures (SAP)–a novel coordination protocol governing component interactions throughout manipulation tasks. Drawing inspiration from Standardized Operating Procedures (SOPs) in human organizations, SAP establishes structured workflows for planning, execution, and verification phases. Our architecture comprises three specialized components: (1) a large reasoning model that decomposes high-level instructions into semantically coherent subgoals, (2) a vision-language-action executor that generates continuous control commands from real-time visual inputs, and (3) a temporal verifier that enables autonomous progression and error recovery through introspective assessment. This SAP-driven closed-loop design supports dynamic self-verification without external supervision. On the LIBERO benchmark, Agentic Robot achieves state-of-the-art performance with an average success rate of 79.6%, outperforming SpatialVLA by 6.1% and OpenVLA by 7.4% on long-horizon tasks. These results demonstrate that SAP-driven coordination between specialized components enhances both performance and interpretability in sequential manipulation, suggesting significant potential for reliable autonomous systems. Project Github: https://agentic-robot.github.io.
@@ 1 Introduction
Inspired by these insights, we design Agentic Robot, a brain-inspired framework that introduces Standardized Action Procedure (SAP)–a novel coordination protocol specifically designed for embodied manipulation tasks. Unlike SOPs, which govern human workflows, SAP encodes the natural cognitive cycle into structured agent interactions for robotic systems. SAP defines the complete agentic loop that governs how our three specialized components–Planner, Executor, and Verifier–coordinate throughout task execution through well-defined interfaces and standardized protocols for information exchange, progress monitoring, and error recovery. Besides, Agentic Robot requires agents to maintain structured interaction protocols throughout the manipulation process. SAP ensures that task decomposition, action execution, and progress verification follow consistent procedures, dramatically reducing error accumulation while enabling robust recovery from failures. More specifically, all components follow strict SAP-defined workflows, ensuring that information handoffs comply with established protocols and eliminating the communication breakdowns that plague existing systems.
@@ 2.1 Overview
We introduce Agentic Robot, an agentic framework that reformulates long-horizon manipulation as a closed perception-reasoning-execution-verification loop, inspired by biological cognition and multi-agent LLM systems [9, 32]. Drawing insights from SOPs that govern effective human workflows, we propose SAP–a novel coordination protocol that structures component interactions throughout the manipulation process. SAP establishes explicit protocols for information exchange, progress monitoring, and error recovery, enabling robust execution of complex manipulation tasks.
@@ 2.3 VLA Executor: Reactive Visuomotor Policy
Despite its stateless design, the executor integrates robust error-handling capabilities through the SAP verification loop. When execution failures occur, the standardized verification mechanism detects issues through visual assessment and triggers specific recovery actions following SAP protocols. If multiple recovery attempts fail, the framework marks the task as failed and halts execution to prevent unsafe behaviors. This closed-loop error detection represents a significant improvement over open-loop methods, enabling autonomous progress through complex manipulation sequences even when individual subgoal executions encounter failures.
@@ 2.4 Verifier: Perception-based Subgoal Assessment and Recovery
We employ Qwen2.5-VL-3B-Instruct [1] as the verifier model to evaluate whether subgoal is complete. The verification prompt follows SAP’s structured format: “Based on the image sequence, has the robot successfully completed [subgoal]?” The model is fine-tuned with LoRA [10] on a dataset of annotated triplets with .
@@
2.5 SAP: Standardized Action Procedures for Coordinated Agentic Control
SAP orchestrates the closed-loop execution within our Agentic Robot framework by structuring interaction across perception, planning, execution, and verification. SAP enables subgoal-level feedback and adaptive recovery, transforming complex long-horizon manipulation into bounded agentic steps. Each SAP cycle at time defines an agentic step as:
where denotes egocentric and third-person views, is the current subgoal, the action taken, and is the verification result. The SAP execution follows four interdependent phases:
1. Multimodal Perception. At each time step, the agent collects:
which provides dual-perspective observations of the workspace.
2. Reactive Execution. The executor translates subgoal into a low-level control signal:
where maps semantic goals and current vision into -DoF actions.
3. Temporal Verification. Every frames (typically 20), the verifier performs a two-stage evaluation:
If , the agent proceeds to the next subgoal. If not, and , a recovery action is triggered.
4. Targeted Recovery. When recovery is needed, a specific correction policy is invoked:
which is followed by re-execution of and repeated verification. After failed recovery attempts, the task is safely terminated.
SAP execution is managed by an asynchronous finite-state machine with component-specific frequencies: the executor operates at ( s), and the verifier at ( s). By enforcing structured control cycles with modular boundaries and layered feedback, SAP enhances agent reliability and interpretability. It supports in-situ correction, isolates errors, and ensures safe recovery—addressing core limitations of open-loop or end-to-end systems in dynamic and uncertain manipulation environments.
@@ 3.1 Experimental Setup
Baselines. We benchmark our approach against the following generalist policies, including state-of-the-art open-sourced models: Diffusion Policy [5], Octo-Base [29], OpenVLA [15], TraceVLA [39], and SpatialVLA [24].
Implementation. Our Agentic Robot consists of three main components: (1) a planner based on GPT-4o [13], which handles visual task decomposition, (2) an executor built on the OpenVLA [15] architecture for vision-language-action alignment, and (3) a verifier using Qwen2.5-VL-3B-Instruct [1], fine-tuned on success and failure episodes. Our recovery strategy raises the gripper to a safe position upon failure detection before re-evaluating. Verification occurs every 20 frames by default.
@@ 3.2 Main Results
| Method | LIBERO-Spatial | LIBERO-Object | LIBERO-Goal | LIBERO-Long | Average | |||||
| SR () | Rank () | SR () | Rank () | SR () | Rank () | SR () | Rank () | SR () | Rank () | |
| Diffusion Policy [5] | 78.3 ± 1.1% | 6 | 92.5 ± 0.7% | 1 | 68.3 ± 1.2% | 6 | 50.5 ± 1.3% | 6 | 72.4 ± 0.7% | 6 |
| Octo-Base (FT) [29] | 78.9 ± 1.0% | 5 | 85.7 ± 0.9% | 5 | 84.6 ± 0.9% | 1 | 51.1 ± 1.3% | 5 | 75.1 ± 0.6% | 4 |
| OpenVLA (FT) [15] | 84.7 ± 0.9% | 3 | 88.4 ± 0.8% | 4 | 79.2 ± 1.0% | 3 | 53.7 ± 1.3% | 4 | 76.5 ± 0.7% | 3 |
| TraceVLA (FT) [39] | 84.6 ± 0.2% | 4 | 85.2 ± 0.4% | 6 | 75.1 ± 0.3% | 5 | 54.1 ± 1.0% | 3 | 74.8 ± 0.5% | 5 |
| SpatialVLA (FT) [24] | 88.2 ± 0.5% | 1 | 89.9 ± 0.7% | 2 | 78.6 ± 0.6% | 4 | 55.5 ± 1.0% | 2 | 78.1 ± 0.7% | 2 |
| Agentic Robot (Ours) | 85.8 ± 0.6% | 2 | 89.0 ± 0.8% | 3 | 81.8 ± 0.8% | 2 | 61.6 ± 1.2% | 1 | 79.6 ± 0.8% | 1 |
@@ 3.3 Long-horizon Manipulation Analysis
We analyze performance on LIBERO-Long, which features multi-step tasks with sequential subgoals. Table 2 reports subgoal-level and overall success rates across 10 manipulation scenarios. Detailed task descriptions are provided in Appendix E.
@@ 3.5 Ablation Study
Verification Quality. Substituting the fine-tuned verifier with a zero-shot VLM yields a substantial drop in success rate to 35.3%, indicating a degradation of 26.5%. This sharp decline suggests that generic models are insufficiently sensitive to subtle changes in scene state, and that domain adaptation is essential for accurate subgoal-level introspection.
Hierarchical structure. Without subgoal decomposition (i.e., vanilla OpenVLA), requiring the system to perform instructions as single, monolithic goals results in a performance drop to 53.7%, a decrease of 8.1%. This supports our main hypothesis that breaking complex tasks into smaller, atomic skills significantly enhances execution reliability and verification accuracy. Each component offers measurable benefits, with fine-tuned verification and hierarchical planning contributing the most substantial improvements. These findings confirm the effectiveness of our Agentic Robot for handling long-horizon manipulation tasks.
@@ 5 Conclusion
This work introduces Agentic Robot, a brain-inspired framework that uses Standardized Action Procedures (SAP) to improve reliability and interpretability in robotic manipulation systems. The framework decomposes complex tasks into coordinated interactions between three specialized components–planner, executor, and verifier–operating through well-defined SAP protocols that mirror biological cognition. By establishing explicit protocols for component communication, progress monitoring, and failure recovery, SAP addresses fundamental limitations in existing manipulation systems while allowing independent component optimization through standardized interfaces. Extensive validation on the LIBERO benchmark demonstrates state-of-the-art performance with 79.6% average success rate, including substantial improvements of 24% on Bowl-Drawer tasks and 21% on Soup-Sauce tasks. This successful integration of brain-inspired architectures with standardized coordination protocols shows how biologically motivated design principles can enhance both performance and interpretability in embodied AI systems.
@@
Appendix A Appendix
The appendix is organized as follows:
Appendix B: We present the detailed experiment setup based on LIBERO.
Appendix C: We present related works for our proposed framework.
Appendix D: We present the atom skill library we used.
Appendix E: We present more detailed results on task division.
Appendix F: We present our proposed framework’s pseudocode.
Appendix G: We present the prompt of LRM for zero-shot task division.
Appendix B Detailed Experiment Setup
B.1 Benchmark
@@ B.2 Baselines
Appendix C Related Works
C.1 Large Reasoning Models
@@
C.2 Vision-Language Models in Robotics
@@
C.3 Vision-Language-Action Models
@@
Appendix D Atom Skill Library
@@
Appendix E Experiments on Task Division
E.1 Description of LIBERO-Long
@@
E.2 Illustration of Task Division
Appendix F Pseudo Code of Agentic Robot

### v2

@@ Abstract
Long-horizon robotic manipulation poses significant challenges for autonomous systems, requiring extended reasoning, precise execution, and robust error recovery across complex sequential tasks. Current approaches, whether based on static planning or end-to-end visuomotor policies, suffer from error accumulation and lack effective verification mechanisms during execution, limiting their reliability in real-world scenarios. We present Agentic Robot, a brain-inspired framework that addresses these limitations through Standardized Action Procedure (SAP)–a novel coordination protocol governing component interactions throughout manipulation tasks. Drawing inspiration from Standardized Operating Procedures (SOPs) in human organizations, SAP establishes structured workflows for planning, execution, and verification phases. Our architecture comprises three specialized components: (1) a large reasoning model that decomposes high-level instructions into semantically coherent subgoals, (2) a vision-language-action executor that generates continuous control commands from real-time visual inputs, and (3) a temporal verifier that enables autonomous progression and error recovery through introspective assessment. This SAP-driven closed-loop design supports dynamic self-verification without external supervision. On the LIBERO benchmark, Agentic Robot achieves state-of-the-art performance with an average success rate of 79.6%, outperforming SpatialVLA by 6.1% and OpenVLA by 7.4% on long-horizon tasks. These results demonstrate that SAP-driven coordination between specialized components enhances both performance and interpretability in sequential manipulation, suggesting significant potential for reliable autonomous systems. Project Github: https://agentic-robot.github.io.
@@ 1 Introduction
Inspired by these insights, we design Agentic Robot, a brain-inspired framework that introduces Standardized Action Procedure (SAP)–a novel coordination protocol specifically designed for embodied manipulation tasks. Unlike SOPs, which govern human workflows, SAP encodes the natural cognitive cycle into structured agent interactions for robotic systems. SAP defines the complete agentic loop that governs how our three specialized components–Planner, Executor, and Verifier–coordinate throughout task execution through well-defined interfaces and standardized protocols for information exchange, progress monitoring, and error recovery. Besides, Agentic Robot requires agents to maintain structured interaction protocols throughout the manipulation process. SAP ensures that task decomposition, action execution, and progress verification follow consistent procedure, dramatically reducing error accumulation while enabling robust recovery from failures. More specifically, all components follow strict SAP-defined workflows, ensuring that information handoffs comply with established protocols and eliminating the communication breakdowns that plague existing systems.
@@ 2.1 Overview
We introduce Agentic Robot, an agentic framework that reformulates long-horizon manipulation as a closed perception-reasoning-execution-verification loop, inspired by biological cognition and multi-agent LLM systems [9, 31]. Drawing insights from SOPs that govern effective human workflows, we propose SAP–a novel coordination protocol that structures component interactions throughout the manipulation process. SAP establishes explicit protocols for information exchange, progress monitoring, and error recovery, enabling robust execution of complex manipulation tasks. Our design is grounded in recent advances across large reasoning models (LRMs), vision-language models (VLMs), and vision-language-action (VLA) systems. We provide a detailed review of these foundations in the Related Works section (see Appendix B).
@@ 2.3 VLA Executor: Reactive Visuomotor Policy
Despite its stateless design, the executor integrates robust error-handling capabilities through the SAP verification loop. When execution failures occur, the standardized verification mechanism detects issues through visual assessment and triggers specific recovery actions following SAP protocols. If multiple recovery attempts fail, the framework marks the task as failed and halts execution to prevent unsafe behaviors. This closed-loop error detection significantly improves system robustness–particularly in long-horizon tasks–by reducing cascading errors. As shown in Section 3.3, our method achieves up to 24% improvement over OpenVLA on the Bowl-Drawer task and 21% on Soup-Sauce, confirming the effectiveness of subgoal-level verification and recovery.
@@ 2.4 Verifier: Perception-based Subgoal Assessment and Recovery
We employ Qwen2.5-VL-3B-Instruct [1] as the verifier model to evaluate whether subgoal is complete. The verification prompt follows SAP’s structured format: “Based on the image sequence, has the robot successfully completed [subgoal]?” The model is fine-tuned with LoRA [10] on a dataset of annotated triplets with . To adapt the verifier to subgoal-level introspection, we fine-tune Qwen2.5-VL-3B-Instruct using LoRA on a compact dataset of approximately 500 annotated triplets. Despite its small scale, the dataset covers diverse subgoal types and visual scenes, and leverages strongly structured prompts to guide learning. This setting demonstrates that even with limited supervision, targeted adaptation can yield effective visual verification in closed-loop execution (see Section 3.5).
@@
2.5 SAP: Standardized Action Procedure for Coordinated Agentic Control
In robotic manipulation, the absence of structured coordination protocols often leads to execution failures, particularly in long-horizon tasks where accumulated errors and lack of systematic verification result in task breakdown. Drawing inspiration from Standardized Operating Procedures (SOPs) that have proven effective in complex collaborative environments, we introduce Standardized Action Procedure (SAP) as a systematic framework that encodes proven coordination patterns into robotic agentic systems.
SAP represents a principled approach to orchestrating closed-loop execution within our agentic robot framework by establishing standardized coordination protocols across perception, planning, execution, and verification components. The core design philosophy rests on three fundamental principles: (1) Modular Decomposition - complex manipulation tasks are systematically decomposed into manageable, verifiable subgoals as illustrated in Fig. 2; (2) Structured Coordination - component interactions follow predefined workflows rather than opportunistic communication; and (3) Adaptive Verification - systematic checkpoints enable early error detection and recovery.
2.5.1 SAP Operational Framework
Each SAP cycle at time constitutes an agentic step that encapsulates the complete perception-planning-execution-verification workflow:
where denotes egocentric and third-person views, represents the current subgoal within the structured task decomposition, is the executed action, and indicates the verification result.
SAP defines four specialized components with standardized interfaces and coordination protocols:
(1) Multimodal Perception. At each time step, the agent collects dual-perspective observations:
which provides comprehensive workspace understanding following standardized observation protocols.
(2) Formulated Plan. The planner converts task instructions into a structured sequence of executable subgoals following standardized decomposition protocols:
where represents the initial visual observation and each subgoal is derived from the Atomic Skill Library.
(3) Reactive Execution. The executor translates subgoal into low-level control signals:
where maps semantic goals and current vision into 7-DoF actions following standardized execution protocols.
(4) Temporal Verification. Every frames (typically 20), the verifier performs systematic evaluation:
where represents the sliding buffer of recent image pairs. If , the agent proceeds to the next subgoal. If not, and , a recovery action is triggered:
SAP execution is managed by an asynchronous finite-state machine with component-specific frequencies: the executor operates at 10 Hz (s), and the verifier at 0.5 Hz (s). By enforcing structured control cycles with modular boundaries and layered feedback, SAP enhances agent reliability and interpretability. It supports in-situ correction, isolates errors, and ensures safe recovery, addressing core limitations of open-loop or end-to-end systems in dynamic and uncertain manipulation environments.
@@ 3.1 Experimental Setup
Baselines. We benchmark our approach against the following generalist policies, including state-of-the-art open-sourced models: Diffusion Policy [5], Octo-Base [28], OpenVLA [14], TraceVLA [38], and SpatialVLA [23]. These methods represent a variety of model paradigms, including diffusion-based control (Diffusion Policy), transformer-based visuomotor policies (Octo-Base), and large-scale vision-language-action models (OpenVLA, TraceVLA, SpatialVLA). For fair comparison, we follow the original hyperparameters and evaluation settings as reported in their respective works without additional tuning. Detailed information are provided in Appendix A.2.
Implementation. Agentic Robot integrates the three modules described in Section 2: a GPT-4o-based planner for subgoal decomposition, an OpenVLA-based executor for visuomotor control, and a fine-tuned Qwen2.5-VL-3B-Instruct verifier for subgoal completion assessment. For error recovery, we raise the gripper to a safe position upon failure detection before re-evaluation. Unless otherwise specified, verification is performed every 20 frames.
@@ 3.2 Main Results
| Method | LIBERO-Spatial | LIBERO-Object | LIBERO-Goal | LIBERO-Long | Average |
| SR (Rank) | SR (Rank) | SR (Rank) | SR (Rank) | SR (Rank) | |
| Diffusion Policy [5] | 78.3 ± 1.1 (6) | 92.5 ± 0.7 (1) | 68.3 ± 1.2 (6) | 50.5 ± 1.3 (6) | 72.4 ± 0.7 (6) |
| Octo-Base (FT) [28] | 78.9 ± 1.0 (5) | 85.7 ± 0.9 (5) | 84.6 ± 0.9 (1) | 51.1 ± 1.3 (5) | 75.1 ± 0.6 (4) |
| OpenVLA (FT) [14] | 84.7 ± 0.9 (3) | 88.4 ± 0.8 (4) | 79.2 ± 1.0 (3) | 53.7 ± 1.3 (4) | 76.5 ± 0.7 (3) |
| TraceVLA (FT) [38] | 84.6 ± 0.2 (4) | 85.2 ± 0.4 (6) | 75.1 ± 0.3 (5) | 54.1 ± 1.0 (3) | 74.8 ± 0.5 (5) |
| SpatialVLA (FT) [23] | 88.2 ± 0.5 (1) | 89.9 ± 0.7 (2) | 78.6 ± 0.6 (4) | 55.5 ± 1.0 (2) | 78.1 ± 0.7 (2) |
| Agentic Robot (Ours) | 85.8 ± 0.6 (2) | 89.0 ± 0.8 (3) | 81.8 ± 0.8 (2) | 61.6 ± 1.2 (1) | 79.6 ± 0.8 (1) |
@@ 3.3 Long-horizon Manipulation Analysis
We analyze performance on LIBERO-Long, which features multi-step tasks with sequential subgoals. Table 2 reports subgoal-level and overall success rates across 10 manipulation scenarios. Detailed task descriptions are provided in Appendix D.
@@ 3.5 Ablation Study
Verification Quality. Substituting the fine-tuned verifier with a zero-shot VLM yields a substantial drop in success rate to 35.3%, indicating a degradation of 26.5%. This sharp decline suggests that generic models are insufficiently sensitive to subtle changes in scene state, and that domain adaptation is essential for accurate subgoal-level introspection. As illustrated in Fig. 5, the zero-shot verifier fails to detect subgoal completion, causing repetitive execution and eventual failure. In contrast, our fine-tuned model enables sequential verification and completion of each subtask. Quantitatively, it improves the LIBERO-Long success rate by +7.8% over the zero-shot baseline, confirming the value of even small-scale domain adaptation.
Hierarchical structure. Without subgoal decomposition (i.e., vanilla OpenVLA), requiring the system to perform instructions as single, monolithic goals results in a performance drop to 53.7%, a decrease of 8.1%. This supports our main hypothesis that breaking complex tasks into smaller, atomic skills significantly enhances execution reliability and verification accuracy. As illustrated in Fig. 6, the lack of subgoal decomposition causes the executor to attempt the full task without the ability to recover from intermediate failures, ultimately skipping necessary steps and leading to task failure. In contrast, our hierarchical agent sequentially verifies and completes each subtask, ensuring successful full-task completion. Each component offers measurable benefits, with fine-tuned verification and hierarchical planning contributing the most substantial improvements. These findings confirm the effectiveness of our Agentic Robot for handling long-horizon manipulation tasks.
@@ 4 Discussion and Limitations
We now reflect on the design of Agentic Robot, highlighting its strengths and pointing out remaining challenges.
@@ 5 Conclusion
This work introduces Agentic Robot, a brain-inspired framework that uses Standardized Action Procedure (SAP) to improve reliability and interpretability in robotic manipulation systems. The framework decomposes complex tasks into coordinated interactions between three specialized components–planner, executor, and verifier–operating through well-defined SAP protocols that mirror biological cognition. By establishing explicit protocols for component communication, progress monitoring, and failure recovery, SAP addresses fundamental limitations in existing manipulation systems while allowing independent component optimization through standardized interfaces. Extensive validation on the LIBERO benchmark demonstrates state-of-the-art performance with 79.6% average success rate, including substantial improvements of 24% on Bowl-Drawer tasks and 21% on Soup-Sauce tasks. This successful integration of brain-inspired architectures with standardized coordination protocols shows how biologically motivated design principles can enhance both performance and interpretability in embodied AI systems.
@@
Appendix A Detailed Experiment Setup
A.1 Benchmark
@@ A.2 Baselines
Appendix B Related Works
B.1 Large Reasoning Models
@@
B.2 Vision-Language Models in Robotics
@@
B.3 Vision-Language-Action Models
@@
Appendix C Atom Skill Library
@@
Appendix D Experiments on Task Division
D.1 Description of LIBERO-Long
@@
D.2 Illustration of Task Division
Appendix E Pseudo Code of Agentic Robot
Appendix F Pseudo Code for Perception-based Verifier